OpenStudy (anonymous):

determine whether u and v are parallel neither or orthogonal u=(-4/3,-5/2,) v=(-8,-15)

OpenStudy (zehanz):

If you calculate the inproduct or dot product of u and v, you can use this to determine the angle between the vectors, with the following formula:\[\cos \theta = \frac{ u.v }{ \left| u \right|.\left| v \right| } \] Hope this helps. ZeHanz